a body of mass 1.5kg is allowed to slide down along a quadrant of a circle from the horizontal position.in reaching to the bottom,its velocity is 8m/s.the work done in overcoming the friction is 12J.the radius of circle is (g=10m/s2)

Grade:11

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
Using the conservation of energy, we have:
mgh=\frac{1}{2}mv^{2}+12
Hence, we get:
1.5\times 10\times h=\frac{1}{2}\times 1.5\times 8^{2}+12\Rightarrow 15h=60\Rightarrow h=4 m
Hence the radius should be equal to 4 m.
Regards
Sumit